OpenStudy (anonymous):

Triangle XYZ where X(7, 2), Y(-3, 8), and Z(-4, 6) is dilated with a scale factor of two-third Where is the dilation located?

OpenStudy (anonymous):

This is my last question!! @mathmale

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Can we assume an Origin of Dilation at (0,0)?

OpenStudy (anonymous):

I guess? I'm really not great with this concept.

OpenStudy (tkhunny):

Calculate the distance of each from the Origin (0,0).

OpenStudy (anonymous):

Would X' be.. \[X'(\frac{ 14 }{ 3},\frac{ 4 }{ 3 })\]

OpenStudy (tkhunny):

Okay, fine. Just do the coordinates. I'm okay with that. Two more to go.

OpenStudy (anonymous):

\[(-2,\frac{ 16 }{ 3 }) and (-\frac{ 8 }{ 3 },4)?\]

OpenStudy (tkhunny):

I think we're done. Good work.

OpenStudy (anonymous):

Thank you!!